OpenStudy (anonymous):

Dividing Polynomials...Please show the steps while answering the questions!!! 1.) (a^2− 28) ÷ (a − 5) 2.) (42x^2-33) ÷ (7x + 7) 3.) (x^2− 74) ÷ (x − 8)

hero (hero):

So for the first one, \(\dfrac{a^2 - 28}{a - 5}\) right?

OpenStudy (anonymous):

yes

hero (hero):

Actually, do you know how to multiply \((a + 5)(a - 5)\) ?

OpenStudy (anonymous):

yes foil method

OpenStudy (anonymous):

i need to know how to do long division with these polynomials

hero (hero):

Or you can apply difference of squares: \((a + b)(a - b) = a^2 - b^2\)

OpenStudy (anonymous):

i kind of know but these are hard with binomials

OpenStudy (anonymous):

the division

OpenStudy (anonymous):

x^2-74 are binomials

hero (hero):

I was showing you a trick method.

OpenStudy (anonymous):

i have to use the division method only,my teacher said.

OpenStudy (anonymous):

help please

hero (hero):

I'll show you how to do the first one.

hero (hero):